### v6

# AudioSolver: Towards Universal Audio Generation by Large-scale Audio Language Models Pre-training

Abstract
Language models (such as GPT-series) have been shown powerful ability in natural language understanding by large-scale generative pre-training strategy. In this study, we formulate all of the audio generation tasks (such as Text-to-speech (TTS), Voice conversion (VC), Speech separation (SS), Text-to-audio (TTA) and Sing voice generation (SVG)) as a sequence modeling problem. We present AudioSolver, a universal audio language model trained by a multi-task learning manner, which faces the following hinders: (1) a proper agent model that can transfer speech, music, sound, and sing data into one shared latent space. (2) a language model that can deal with long sequences. To solve these problems, we first train a universal audio codec model in large-scale audio-only datasets. Then, we propose a multiscale GPT structure to solve the long sequence problem in audio generation. We scale the training dataset to 150k hours and the model size into 1B. We demonstrate that our universal audio language model has strong generalization ability, which got the SOTA performance in many zero-shot tasks. Furthermore, we find that the pre-trained audio language model can be easy to fine-tune in a new task. To facilitate the research in audio language models, we release AudioSolver, the first open-source audio language model tools, which empowers humans to train audio language models with unprecedented ease.
1 Introduction
Large language models (LLMs), brown2020language; chowdhery2022palm; OpenAI have been shown powerful language understanding ability, which can solve many natural language processing (NLP) related tasks. Large-scale generative pre-training strategy and increased model parameters are two important factor for the sucess of LLMs. Inspired by the success of LLMs, audio language models got a lot of attention in the context of audio generation. AudioLM borsos2023audiolm is one of the representative work. AudioLM introduced a hierarchical approach which combines two types of audio tokens, with high-level semantic tokens extracted from W2V-BERT chung2021w2v being used to condition the generation of low-level acoustic codes from a neural audio codec model zeghidour2021soundstream. Following AudioLM, a lot of similar works have been proposed to solve audio generation tasks, e.g. VALL-E wang2023neural and SPEARTTS kharitonov2023speak proposed to solve text-to-speech (TTS) task; MusicLM agostinelli2023musiclm and MusicGen copet2023simple proposed to solve text-to-music task; AudioGen kreuk2022audiogen proposed to solve text-to-sound task; Make-A-Voice huang2023make proposed to solve voice conversion tasks. Although these models can get amazing performance for special tasks, their target is far from the LLMs, that is, using a unified model to solve all of tasks. Furthermore, to reserve the acoustic information that influence the user’s auditory sense in the audio, residual vector quantization (RVQ) technique zeghidour2021soundstream is used. RVQ reduces the information loss through increasing the VQ layer, which also increases the burden for generation models (we refer it as long sequence modeling problem). To avoid modelling a very long sequence, VALL-E and AudioLM are both proposed to training a multi-stage models. SPEAR-TTS and MusicGen train a one single stage model, but such models require higher GPU memory. In our practice, training a GPT-like models to deal with long sequence (more than 2000 tokens) is unstable. Thus, long sequence modelling problem is a big hinder in audio language model fields.
Table 1: The audio generation tasks supported by UniAudio and prior works. * means the prior works that do not claim to support that task but we believe they have the potential to. TTS: text-to-speech; VC: voice conversion; SE: speech enhancement; TSE: target speech extraction; SVS: singing voice synthesis; Sound: text-to-sound generation; Music: text-to-music generation; A-Edit: Audio Edit; SD: speech dereverberation; I-TTS: Instructed TTS; S-Edit: Speech edit.
Model TTS VC SE TSE SVS Sound Music A-Edit SD I-TTS S-Edit
YourTTS (yourtts)
VALL-E (valle)
MusicLM (valle)
SPEARTTS (speartts)
NaturalSpeech2 (naturalspeech2)
Make-A-Voice (makeavoice)
Maga-TTS (megatts)
VoiceBox (voicebox)
AudioLDM2 (audioldm2)
SpeechX (speechx)
UniAudio (ours)
In this study, we adopt the idea of large-scale multi-task pre-training from GPT-series radford2018improving; radford2019language; brown2020language. We formulate all of audio related generation tasks as a unified audio language model problem, where we meet with the following research problems: (1) how to choose a proper latent space that different audio data (such as speech, sound, music and sing) can be mapping into the shared discrete space; (2) how to solve the long sequence modeling problem when we want to train a one stage model while keeping more acoustic details; (3) whether large-scale multi-task training can bring improvement; (4) whether pre-trained models can be used for new tasks. In this paper, we will answer these questions. The contributions of this work are:
  • $\bullet$
    We propose a universal audio generation framework to solve all of the audio related generation tasks. To our best knowledge, this is the first work to consider using a language model to solve all of the audio generation tasks.
  • $\bullet$
    To solve the long sequence modelling problem in audio generation, we propose a multiscale GPT structure, that using a global GPT to handle the semantic information modeling, and a local GPT to handle the fine-grained acoustic information modeling. By using our multiscale GPT structure, more acoustic details can be modeling without introducing long sequence.
  • $\bullet$
    Our universal audio language models perform strongly generalization ability. We evaluate our model on zero-shot TTS, zero-shot voice converison, zero-shot speech enhancement, zero-shot speaker extraction, zero-shot sing voice generation, zero-shot text-to-sound and music generation tasks. Experimental results show that our proposed method got the SOTA performance.
  • $\bullet$
    Our pre-trained models can be easy to finue-tune on new tasks. We fine-tune our pre-trained models on speech edit, audio edit, speech dereverberation and instruct-controllable expressive TTS (InstructTTS for short) tasks.
2 AudioSolver
In this section, we introduce the details of AudioSolver. We first introduce how to formulate all of the generation tasks as the sequence modelling problem. Then we give the details of our proposed multi-scale GPT structure to deal with any length sequence. The details of AudioSolver as Figure 1 shows.
2.1 Universal Input Representation
AudioSolver is a universal audio language model that can perceive general modalities, follow instructions, learn in context, and generate outputs. The context can be the phone-sequence, task label, instruction, audio sequence and so on. Based on the previous context, the model learns to generate audio in an auto-regressive manner. Our AudioSolver’s backbone is a Transformer-based causal language model. Following GPT-1, we use the special token to represent different task. We classify the context input into two groups: the first group is discrete sequence, such as phone sequence, audio sequence, task label. The second group is the continuous embedding representations, such as the text representations extracted by a T5 encoder (raffel2020exploring). For these discrete sequence, we directly define a learnable embedding table for them, the embedding table can be optimized with the network. For these continuous embedding, we use a linear layer map them into the same dimension as the learnable embedding dimension, then combine with other embedding sequence. Figure 1 (a) shows six tokenizers to help us realize universal input representation. LabelTokenizer is defined by ourselves, in this study, we define the first 128 tokens in the vocabulary as special tokens, which supports users to define any task label. The later token will start from 128. For PhoneTokenizer and MIDITokenizer, we just follow the map dict that transfer the symbol sequence into token sequence. For SemanticTokenizer, we use the pre-trained HuBERT model and a k-means cluster to quantize the audio signal into semantic token sequence. The HuBERT model’s downsample rate is 320, which means that 1 second audio signal with 16khz sample rate will produce 50 tokens. For AudioTokenzier, we use the neural audio codec models to compress the audio signal into discrete representations. The details of audio codec models will be introduced in the next section.
Figure 1: The overview of AudioSolver.
2.2 Universal Neural Audio Codec
In this study, audio codec models play an important role, which can compress any type audio data (speech, sound, music, sing) into a shared discrete latent space. By using the codec models, we can use the audio language models to generate any type audio. In general, a neural audio codec model consists of three part: (1) an encoder $E$ that takes the audio signal $\bm{x}$ and generates a latent feature representation $\bm{z}$; (2) several residual vector quantization (RVQ) layers that produces a compressed representation $\bm{z}_{q}$; and a decoder that recovers the audio signal $\bm{\hat{x}}$ from the compressed latent representation $\bm{z}_{q}$. In practice, each quantization layer will maintain a codebook (a group of learnable embedding vector). Thus $\bm{z}_{q}$ can be denoted by the index of embedding vector. Assume the down-sample times is $S=320$ in the encoder, and the number of VQ layers $n_{q}=3$, so that 1 second 16khz audio data can be represented by $50\times 3$ tokens. Although Encodec have release the pre-trained models, Encodec results in poor audio quality, which has been verified in VALL-E. HiFi-Codec can bring better audio quality, but it only be trained in speech data.
Table 2: Performance comparison between different universal audio codec models.
Model $n_{q}$ PESQ STOI PESQ STOI PESQ STOI PESQ STOI PESQ STOI
Type - Speech Sound Music Sing AVE
Encodec 8 2.18 0.79 2.23 0.48 1.86 0.57 1.95 0.76 2.05 0.649
I-Encodec 8 3.36 0.878 2.67 0.54 2.31 0.65 3.49 0.89 2.95 0.739
I-Encodec 4 3.11 0.86 2.5 0.51 2.08 0.59 3.27 0.86 2.73 0.71
I-Encodec 3 2.96 0.85 2.42 0.49 1.99 0.57 3.13 0.85 2.62 0.69
HiFiCodec 4 3.21 0.875 2.58 0.53 1.9 0.57 3.5 0.87 2.79 0.71
In this study, we first train a universal audio codec models. We explore two types audio codec models: (1) Improved Encodec model (I-Encodec), we use the same encoder-decoder backbone from open source audio codec model  [^1] https://github.com/facebookresearch/encodec, combined with our improved multi-scale frequency domain discriminator (refer to the Appendix); (2) universal HiFi-Codec model, which we follow the official open source code  [^2] https://github.com/yangdongchao/AcademiCodec, the difference is that we use more audio data. Table 2 shows the reconstruction performance comparison between Encodec and our trained codec models. We can see that our codec models are better than Encodec. Furthermore, using more VQ layers can bring better reconstruction performance. Theoretically, the limit of audio generation model is the reconstruction performance. However, using more VQ layers also bring the burden for the generation models. In the following, unless specifically stated, we follow the setting of SPEARTTS, that using I-Encodec with 3 VQ layers.
Table 3: The comparison between different model structure.
Structure Codec MOS ($\uparrow$) Max audio duration Min GPU memory Training Stage
AudioSolver I-Codec-3 - 30s 16G 1
AudioSolver I-Codec-4 0.64 30s 16G 1
AudioSolver I-Codec-8 0.64 30s 16G 1
VALL-E Encodec-8 - 20s 32G 2
SPEARTTS I-Codec-3 - 10s 64G 1
SPEARTTS I-Codec-8 N/A 1s 64G 1
2.3 Multi-scale GPT Model
Long sequence modeling problem (due to the hierarchical representations in audio codec) is one of the core problems in audio language models. Previous works try to solve this by training multi-stage models, e.g. AudioLM tries to train three separated models, VALL-E trains two independent models. SPEARTTS and MusicGen both find that flattening all codebooks into one sequence brings the best performance. The drawback is that such a pattern will result in long sequences and cost large-scale GPU resources. To combine the advantage of flattening patterns and reducing the sequence length, we propose a multiple-scale GPT structure inspired by yu2023megabyte. Our motivation is that the tokens belonging to different codebooks should enjoy the same semantic information along the same time. Specifically, assuming the audio codec sequence is $a_{1,1},a_{1,2},a_{1,3},...,a_{n,1},a_{n,2},a_{n,3}$ after the flattening operation, where $a_{i,1},a_{i,2},a_{i,3}$ should enjoy the same semantic guidance $s_{i}$. Based on this assumption, we first use a global GPT to model the semantic information (the sequence length is $n$, not previous $3\times n$), then we use a local GPT to generate $a_{i,1},a_{i,2},a_{i,3}$ (the sequence length is 3). By using our model, the max sequence length is $n$, which effectively removes the long sequence problem caused by multi-codebooks. Figure 1 (b) presents the overview of the multi-scale GPT model, which consists of three parts: (1) Patch embedder that combine multi-sequence as one patch (we directly use the concatenate operation along the feature dimension); (2) Global GPT module that using patched features to model the semantic-level information; (3) Local GPT module that using the semantic-level features to model acoustic details. To make the local GPT can parallel deal with all of patch, we repeat other types of tokens $n_{q}$ times, so that the semantic information can be align very well.
2.3.1 Patch Embedder
Assuming the input sequence is $\{x_{i}\}_{i=0}^{i=T-1}$. Patch Embedder aims to reshape the sequence into a patch sequence with the length of $K=\frac{T}{P}$, where $P$ denotes the patch size. Firstly, a global embedding table $E_{g}\in\mathbb{R}^{V\times D_{g}}$ is used to map the input sequence into $\{\bm{x}_{i}\}_{i=0}^{i=T-1}$, where $\bm{x}_{i}\in\mathbb{R}^{D_{g}}$. The positional embedding is used, thus the hidden embedding can be $\bm{h}_{i}^{embed}=\bm{x}_{i}+\bm{p}_{i}$. The Patch Embedder aims to reduce the sequence length of $\bm{H}^{embed}$ by patch concatenate:
$\bm{H}^{embed}\in\mathbf{R}^{T\times D_{g}}\rightarrow\bm{H}^{global-in}\in\mathbb{R}^{K\times(P\cdot D_{g})}$ (1)
2.3.2 Global GPT and Local GPT
After we get the patched features, we can directly feed these features into the global GPT models (a decoder-only Transformer model).
$\bm{H}^{global-out}=GPT^{global}(\bm{H}^{global-in})\in\mathbb{R}^{K\times(P\cdot D_{g})}$ (2)
Then we reshape $\bm{H}^{global-out}$ into the shape of $K\times P\times D_{g}$. Furthermore, we combine $\bm{H}^{global-out}$ with the local embedding (by a local embedding table $E_{l}\in\mathbb{R}^{V\times D_{l}}$):
$\bm{H}^{local-in}=\bm{W}_{g}\bm{H}^{global-out}+\bm{H}^{embed-local}$ (3)
where $\bm{W}_{g}$ denotes the linear mapping layer. Similarly, the local GPT is also a decoder-only Transformer.
$\bm{H}^{local-out}=GPT^{local}(\bm{H}^{local-in})\in\mathbb{R}^{K\times P\cdot D_{l}}$ (4)
Finally, we can compute the probability distribution over the vocabulary at each position based on $\bm{H}^{local-out}$. Lastly, we can calculate the Cross Entropy loss with the ground truth label.
3 Experiments
4 Conclusion and Future works
In this study, we present AudioSolver, the first universal audio language model that can solve any audio generation tasks. A lot of experiments have confirmed the effectiveness of AudioSolver.
Appendices
Appendix A Experimental Setup
This appendix describes experimental setups in detail, including data statistics, model architecture and optimization strategy.
A.1 Data Description
12 public datasets are adopted in this work for training. Besides, several test sets are additionally used only for zero-shot evaluation. The statistics of these datasets are in Table 4. Datasets adoption for each task is described in Table 5. Note some datasets are adopted by more than one task.
Table 4: Data statistics
Dataset Type Annotation Volume (hrs)
Training
LibriLight (librilight) speech - 60k
LibriTTS (libritts) speech text 1k
MLS (pratap2020mls) speech - 20k
AudioSet (audioset) sound - 5.8k
AudioCaps (audiocaps) sound text description 500
WavCaps (mei2023wavcaps) sound text description 7k
Million Song Dataset (msd) music text description 7k
OpenCPOP (opencpop) singing text, MIDI 5.2
OpenSinger (opensinger) singing text, MIDI 50
AISHELL3 (shi2020aishell) speech text 85
PromptSpeech (guo2023prompttts) speech text, instruction 200
openSLR26,openSLR28 (ko2017study) room impulse response - 100
Test
LibriSpeech test-clean librispeech speech text 8
VCTK (vctk) speech text 50
TUT2017 Task1 (tut2017) Noise - 10
Cloth (drossos2020clotho) Sound text description 3
MusicCaps (musiclm) Music text description 15
M4Singer(m4singer) singing text, MIDI 1
Table 5: Dataset adoption of all tasks
Task Training dataset Test set Train Volume (hrs)
Training Stage
TTS Librilight LibriSpeech clean-test 60k
VC Librilight VCTK 60k
SE MLS, Audioset TUT2017 Task1, VCTK 20k
TSE MLS Libri2Mix test set 10k
Sound AudioCaps, WavCaps Cloth test set 7k
Music MSD MusicCaps 7k
Singing OpenCPOP, OPenSinger, AISHEELL-3 M4Singer test set 150
Fine-Tuning Stage
I-TTS PromptSpeech PromptSpeech test set 200
Speech dereverberation LibriTTS, openSLR26, openSLR28 LibriTTS test set 100
Speech edit LibriTTS LibriTTS test set 100
Audio edit AudioCaps, WavCaps AudioCaps test set 500
Sum - - 166k
A.2 Model Configuration
The model configuration of the proposed multi-scale Transformer is described in Table 6.
Table 6: Model configuration (with $n_{q}=3$)
Hyper-parameter Global Transforemr Local Transformer
#layer 24 8
#Attention dim 1536 1536
#Attention head 12 12
#Feed-Forward dim 6144 6144
#Params (M) 744 238
Max context length (in #tokens) 3,000 3
Causality Yes Yes
A.3 Optimization
The optimization configurations adopted in both the training and fine-tuning stages are presented in Table 7
Table 7: Optimization Configuration
Hyper-parameter Pre-training Fine-Tuning
Batch Size (#patches/GPU) 8k 8k
Peak Learning Rate 1e-4 1e-5
Warm-up Steps 10000 1000
Training Steps 800k 50k
Learning rate decay Noam (vaswani2017attention) Noam (vaswani2017attention)
Appendix B The Details of Experiments
This section presents detailed experimental results on each task. In the following, if the training set and test sets come from different datasets, we label them as zero-shot settings.
B.1 TTS and VC tasks
For TTS tasks, UniAudio is compared with the many previous SOTA models, Table 8 presents the results. For FastSpeech 2, we only conduct QMOS evaluation as its implementation adopts speaker id as input  [^3] https://github.com/ming024/FastSpeech2. We can see that UniAudio obtains better performance in terms of WER, SIM than YourTTS, VALL-E, NaturalSpeech 2 and Make-A-Voice. Compared with VoiceBox, UniAudio also gets comparable performance in terms of objective metrics. From the MOS evaluation, we can see that UniAudio can generate high-quality speech compared with previous SOTA works. Furthermore, UniAudio realizes the best zero-shot clone ability (e.g. SMOS is 3.56 and SIM is 0.708). More experiments, such as cross-lingual zero-shot TTS and Mandarin Chinese speech synthesis can be found in demo page. For VC task, we conducted experiments on VCTK dataset, we randomly chose 200 audio pairs. PPG-VC and YourTTS are trained on small-scale datasets. Make-A-Voice and LM-VC  [^4] We seek help from the authors, they provide the inference results. are trained on large-scale datasets as the same as UniAudio. Compared with previous work, UniAudio got better performance in voice conversion tasks.
Table 8: The performance comparison with previous SOTA methods in TTS and VC tasks. We do not conduct MOS evaluation for VALL-E, SPEARTTS and VoiceBox due to the models are not released.
Model Zero-shot SIM ($\uparrow$) WER ($\downarrow$) MOS ($\uparrow$) SMOS ($\uparrow$)
Text-to-Speech
GroundTruth - - 1.9 3.99$\pm$0.08 -
FastSpeech 2 (fastspeech2) - - 3.81$\pm$0.10 -
YourTTS (yourtts) 0.337 7.7 3.66$\pm$0.07 3.02$\pm$0.07
VALL-E (valle) 0.580 5.9 - -
Make-A-Voice (TTS) (makeavoice) 0.498 5.7 3.74$\pm$0.08 3.11$\pm$0.06
NaturalSpeech 2 (naturalspeech2) 0.620 2.3 3.83$\pm$0.10 3.11$\pm$0.10
SPEAR-TTS (speartts) 0.560 / - -
VoiceBox (voicebox) 0.681 1.9 - -
UniAudio 0.708 2.0 3.81$\pm$0.07 3.56$\pm$0.10
Voice Conversion
GroundTruth - - 3.25 3.74$\pm$0.08 -
PPG-VC (ppg-vc) 0.78 12.3 3.41$\pm$0.10 3.47$\pm$0.10
YourTTS (yourtts) 0.719 10.1 3.61$\pm$0.10 3.26$\pm$0.10
Make-A-Voice (VC) (makeavoice) 0.678 6.2 3.43$\pm$0.09 3.47$\pm$0.10
LM-VC (lm-vc) 0.820 4.91 3.41$\pm$0.08 3.17$\pm$0.09
UniAudio 0.868 4.8 3.54$\pm$0.07 3.56$\pm$0.07
B.2 Speech Enhancement and Target Speaker Extraction
For the SE task, we compare with previous SOTA methods, including discriminative methods (such as FullSubNet and FullSubNet+) and generative methods (such as SGMSE+ and NADiffuSE). Note that the CDiffuSE and NADiffuSE are both trained on the voicebank-demand dataset. Other models never saw the VCTK dataset in the training stage. We obtain the inference results based on their open-source models. Table 9 presents the results, we can see that UniAuido obtains the best DNSMOS score. The PESQ and VISQOL scores are lower than other SOTA methods, we think these metrics may not accurately assess the performance of generative methods. The similar finding is also observed in previous literature (tokensplit) that the signal-level evaluation metrics may not be suitable for generative methods. In contrast, we recommend using DNSMOS and MOS scores as the main metrics. UniAuido can get good results in extremely noisy environments, we recommend readers refer to the demo page. For the TSE task, we conducted experiments on the LibriMix test set. The popular TSE systems: VoiceFilter  [^5] https://github.com/Edresson/VoiceSplit and SpeakBeam [^6] https://github.com/BUTSpeechFIT/speakerbeam are used as baseline systems. As Table 9 shows, we can see that UniAudio obtains the best performance in terms of DNSMOS and MOS.
Table 9: The performance of SE and TSE tasks comparison with previous SOTA methods.
Model Zero-shot PESQ ($\uparrow$) VISQOL($\uparrow$) DNSMOS($\uparrow$) MOS($\uparrow$)
Speech Enhancement
CDiffuSE (lu2022conditional) 1.88 1.21 2.54 -
NADiffuSE (wang2023nadiffuse) 2.96 2.41 3.03 3.30$\pm$0.08
SGMSE+ (richter2023speech) 3.21 2.72 3.29 3.56$\pm$0.08
FullSubNet (hao2021fullsubnet) 3.21 2.77 3.37 3.61$\pm$0.10
FullSubNet+ (chen2022fullsubnet+) 3.41 2.99 3.34 3.42$\pm$0.08
UniAudio 2.63 2.44 3.66 3.68$\pm$0.07
Target Speaker Extraction
SpeakerBeam (vzmolikova2019speakerbeam) 2.89 2.25 3.18 3.68$\pm$0.1
VoiceFilter (wang2018voicefilter) 2.41 2.36 3.35 3.43$\pm$0.09
UniAudio 1.88 1.68 3.96 3.72$\pm$0.06
B.3 Singing Voice Synthesis
Following Make-A-Voice, we conduct experiments on the M4Singer test set. We compare the generated singing samples with other systems, including 1) Diffsinger; 2) Make-A-Voice, a two-stage audio language model for singing voice generation. As illustrated in Table 10, we can see that UniAudio gets comparable results with Make-A-Voice and Diffsinger.
Table 10: Quality and style similarity of generated samples in singing voice synthesis.
Model MOS ($\uparrow$) SMOS ($\uparrow$)
Diffsinger (liu2022diffsinger) 3.94$\pm$0.02 4.05$\pm$0.06
Make-A-Voice (makeavoice) 3.96$\pm$0.03 4.04$\pm$0.05
UniAudio 4.08$\pm$0.04 4.04$\pm$0.05
B.4 Text-to-sound and text-to-music generation
The text-to-sound generation task has attracted great interest in audio research. Following Diffsound (diffsound), most of the methods evaluate their systems on the AudioCaps (audiocaps) test set. However, we found that if the training data includes the AudioCaps data, the model is easy to overfit with AudioCaps. As a result, the best performance can be obtained when the model only trains on the Audiocaps. In this study, we conduct a zero-shot evaluation on the Cloth test set (drossos2020clotho). Table 11 shows the results. We can see that UniAudio obtains better performance than Diffsound and AudioLDM. Compared to recent SOTA models, such as Tango and Make-an-Audio 2, UniAudio also gets comparable performance. For the text-to-music task, we follow MusicGen (musicgen), evaluating our methods on MusicCaps (musiclm). Compared with previous SOTAs, UniAudio gets a comparable performance with other models. From the MOS evaluation performance, we can see that MusicGen is better than our current models. We speculate one of the reasons is that MusicGen uses a large-scale high-quality dataset (20k hours).
Table 11: Text-to-sound and text-to-music evaluation. We report the subjective metrics including FAD($\downarrow$), and KL($\downarrow$). Furthermore, we also conduct objective evaluation. Note that the training data of AudioGen includes Cloth datatset, thus can not be seen as zero-shot setting.
Model Training Data (Hours) FAD KL OVL. REL.
Text-to-Sound Generation
Reference / / / 70.47$\pm$1.9 78.84$\pm$1.5
Diffsound 2k 7.8 6.53 - -
AudioGen 4k 2.55 2.5 63.84$\pm$2.1 72.12$\pm$1.8
Tango 3.3k 3.61 2.59 66.2$\pm$1.7 68.57$\pm$1.5
Make-an-Audio 2 8.7k 2.13 2.49 61.52$\pm$1.6 69.9$\pm$1.5
AudioLMD 9k 4.93 2.6 60.95$\pm$1.9 65.7$\pm$1.8
UniAudio 7k 3.12 2.57 61.9$\pm$1.9 66.1$\pm$1.5
Text-to-Music Generation
Riffusion - 14.8 2.06 - -
Mousai - 7.5 1.59 - -
MusicLM 280k 4.0 - - -
Noise2Music 280k 2.1 - - -
MusicGen 20k 4.52 1.41 73.28$\pm$1.5 71.28$\pm$1.7
UniAudio 8k 3.65 1.87 67.85$\pm$1.70 70.0$\pm$1.5
B.5 Audio Edit
Audio edit aims to edit the original audio based on Human’s instruction. AUDIT (wang2023audit) is the SOTA model in audio edit task, which designs a data simulation strategy to get triplet training and test data (e.g., {audio, audio, text}). The authors set 5 different tasks, including adding, dropping, replacing, inpainting and super-resolution, and simulated large-scale data for each task. To validate that our pre-trained model can be fine-tuned with small-scale data, we choose adding, dropping and super-resolution tasks to fine-tune simultaneously. To finish the fine-tuning process, we define a new task label: Audit_task. The experimental results as Table 12 shows. We can observe that: (1) UniAudio can get better performance with the previous SOTA model. (2) Fine-tuning pre-trained UniAudio can get better performance than training it from scratch, which further validates the effectiveness of pre-training a model on large-scale training data.
Table 12: Audio edit task evaluation.
Type Model FD KL
Adding task
AUDIT 21.80 0.92
UniAudio (scratch) 20.2 0.99
UniAudio (fine-tune) 19.69 0.934
Dropping task
AUDIT 22.40 0.95
UniAudio (scratch) 27.76 1.38
UniAudio (fine-tune) 23.1 1.10
Super-Resolution task
AUDIT 18.14 0.73
UniAudio (scratch) 11.51 0.29
UniAudio (fine-tune) 10.54 0.289
Table 13: Quality and style similarity of generated samples for Instructed TTS task.
Model MOS ($\uparrow$) SMOS ($\uparrow$)
GT 3.77$\pm$0.07 3.85$\pm$0.08
UniAudio (scratch) 3.62$\pm$0.07 3.67$\pm$0.08
UniAudio (tuning) 3.61$\pm$0.09 3.71$\pm$0.09
B.6 Instructed TTS
Using instruction to guide speech synthesis has received great attention (guo2023prompttts; yang2023instructtts). In this part, we fine-tune the UniAudio model on the PromptSpeech (guo2023prompttts) dataset. Furthermore, we also try to train a UniAudio model from scratch with the PromptSpeech dataset. Different from previous works that designed special style encoders to capture the style information from text descriptions, we directly use the T5 text encoder to extract representations from text and then combine it with the phoneme sequence input to the UniAudio, which is more convenient. [^7] Note that the authors of PromptTTS (guo2023prompttts) told us their objective metrics tools, checkpoints, and generated samples have been lost due to the machine errors. Thus we cannot fairly compare with them. Table 13 shows the results, we can see that UniAudio has good performance in terms of style control and speech quality when compared with the ground truth samples.
B.7 Speech Dereverberation
For the speech dereverberation task, we use the Room Impulse Response (RIR) data from the openSLR26 and openSLR28 dataset, and the speech data from the LibriTTS clean part. We simulate about 100 hours of training data and 1 hour of test data. We compare with previous SOTA systems, such as FullSubNet, FullSubNet+ and SGMSE+. Table 14 presents the results. We can see that UniAudio obtains the SOTA performance in speech dereverberation tasks with small-scale training data in terms of DNSMOS metric. Similar with speech enhancement task, we speculate that PESQ may not suitable for the generative methods.
Table 14: Results comparison with previous speech Dereverberation systems.
Model PESQ ($\uparrow$) DNSMOS($\uparrow$)
SGMSE+ 2.87 3.42
FullSubNet 2.29 3.32
FullSubNet+ 2.27 3.25
UniAudio (scratch) 1.23 3.18
UniAudio (tuning) 2.13 3.51
B.8 Speech Edit
For the speech edit task, we use the LibriTTS dataset. In practice, we randomly choose some words to mask in the training stage. We expect the model to recover the whole speech based on the phoneme sequence. In the inference stage, we can mask the region that we want to update in the speech and input the new words so that the model can edit the speech. For this task, we take the TTS system that regenerates a complete waveform from the whole sentence to be edited as the baseline. In the evaluation, we mainly validate three situations: (1) word replacement; (2) insert a new word; and (3) delete a word. For each situation, we randomly chose 10 sentences from the LibriTTS test clean set.
Appendix C Ablation study
C.1 The influence of multi-task training
In this part, we explore whether multi-task training can bring better performance than task-specific training. To answer this question, we use the same model trained on different tasks, respectively. Table 15 shows the experimental results, UniAudio (single) means that the model is trained on a single task. We observe that multi-task training brings the gain over all of the tasks. In Appendix D, we give some potential reasons why multi-task training can bring improvement.
Table 15: The ablation study of the effectiveness of multi-task training.
Task Model Objective Evaluation Subjective Evaluation
Metrics Results Metrics Results
[…]
/ SMOS(($\uparrow$( ↑) 3.77$\pm$0.06 / 3.46$\pm$0.10
UniAudio 0.71 / 2.0 3.81$\pm$0.07 / 3.56$\pm$0.10
[…]
/ SMOS(($\uparrow$( ↑) 3.45$\pm$0.07 / 3.44$\pm$0.07
UniAudio 0.87 / 4.8 3.54$\pm$0.07 / 3.56$\pm$0.07
[…]
/ VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 2.35 / 2.30 / 3.45 MOS(($\uparrow$( ↑) 3.65$\pm$0.08
UniAudio 2.63 / 2.44 / 3.66 3.68$\pm$0.07
[…]
/ VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 1.97 / 1.61 / 3.93 MOS(($\uparrow$( ↑) 3.58$\pm$0.08
UniAudio 1.88 / 1.68 / 3.96 3.72$\pm$0.06
[…]
/ SMOS(($\uparrow$( ↑) 4.14$\pm$0.07 / 4.02$\pm$0.02
UniAudio 4.08$\pm$0.04 / 4.04$\pm$0.05
[…]
/ REL ()($\uparrow$)( ↑ ) 60.0$\pm$2.1 / 61.2$\pm$1.8
UniAudio 3.12 / 2.6 61.9$\pm$1.9 / 66.1$\pm$1.5
[…]
/ REL ()($\uparrow$)( ↑ ) 64.4$\pm$2.1 / 66.2$\pm$2.4
UniAudio 3.65 / 1.9 67.9$\pm$1.7 / 70.0$\pm$1.5
[…]
UniAudio 17.78 / 0.77 -
[…]
UniAudio 2.13 / 3.51 -
[…]
UniAudio - 3.61$\pm$0.09 / 3.71$\pm$0.09
[…]
UniAudio 5.12 3.82$\pm$0.06
C.2 Fine-tuning the pre-trained model on the new task will influence the performance on previous tasks?
In this part, we conduct experiments to explore whether fine-tuning the pre-trained model on new tasks will influence the performance of previous tasks. We evaluate the pre-trained UniAudio model (trained on 7 tasks) and fine-tuned UniAudio model (fine-tuned on 4 new tasks) on 7 tasks. Figure 2 shows the results. We can see that the performance does not significantly drop on previous training tasks, which demonstrates that UniAudio has the potential to add new tasks continuously without losing previous task knowledge.
Figure 2: Performance comparison over 7 audio generation tasks before/after fine-tuning.
C.3 The influence of data quantity
In this part, we conduct experiments to explore the influence of data quantity, we give three settings: (1) using all of the data; (2) using $1/2$ training data for each task; (3) using $1/4$ training data for each task. We present the results in Figure 3. Based on the experimental results, this work claims that the data quantity is a key point to building a strong audio foundation model. In the future, we will explore to use of more unlabeled data to help improve the performance.
Figure 3: Performance comparison over different data quantity.
Appendix D Why UniAudio Can Work Well?
From the previous discussions, we can see that the universal modeling strategy brings improvement for different tasks. In this part, we try to give some potential explanations.
(1) Deterministic latent space: we formulate different modalities into a deterministic latent space (fixed vocabulary) by tokenization. Different tokens can be seen as specific ’words’, and we can use a next-token prediction strategy to train the model. Similar to GPT-series (radford2018improving; radford2019language), such strategy creates the opportunity for the model to learn the intrinsic properties of audio and the interrelationship between audio and other modalities.
(2) Shared information between different types of audio: Although multiple types of audio (speech, sounds, music, and singing) present significant differences in the time domain or frequency domain, neural audio codec models effectively capture their shared information (rethinking the working principle of neural codecs, which similar information will be allocated the same token id). Due to the shared information that exists in different types of audio, multi-task training can be seen as increasing training data for each task.
(3) Data augmentation perspective: We speculate that multi-task training can be viewed as data augmentation for some tasks. Considering the TTS and VC task’s definition:
TTS: <phoneme_sequence> <prompt> <audio_sequence>
VC: <semantic_token> <prompt> <audio_sequence>
We can see that the difference in task formulation for TTS and VC is that they use different ways to denote the phonetic information. In essence, they carry the same phonetic information. The difference is that semantic tokens include the duration information. Thus we can view the phoneme sequence as a special semantic sequence that drops the duration information. Such dropping operation is widely used as a data augmentation strategy (specaugment).
Appendix E The details of Audio Codec Models
[…]
Type Speech (VCTK) Sound (cloth) Music (musiccaps) Sing (m4sing) Average
(vctk) (drossos2020clotho) (musiclm) (m4singer) -
Model $n_{q}$ FPS TPS PESQ STOI PESQ STOI PESQ STOI PESQ STOI PESQ STOI
Encodec 8 75 600 2.18 0.79 2.23 0.48 1.86 0.57 1.95 0.76 2.05 0.65
Ours 3 50 150 2.96 0.85 2.42 0.49 1.99 0.57 3.13 0.85 2.62 0.69
Ours 4 50 200 3.11 0.86 2.5 0.51 2.08 0.59 3.27 0.86 2.73 0.71
Ours 8 50 400 3.36 0.88 2.67 0.54 2.31 0.65 3.49 0.89 2.95 0.74
In this part, we give more details about our neural audio codec model in Section LABEL:sec:audio. We adopt a similar encoder-decoder framework with the Encodec model, the difference includes: (1) we replace the multi-scale STFT-based (MS-STFT) discriminator as our multi-scale Mel-based discriminator. (2) We rewrite the vector quantization implementation  [^8] Please refer to our source code to find the details. based on Encodec’s open-source version  [^9] https://github.com/facebookresearch/encodec/blob/main/encodec/quantization/core_vq.py, making it more suitable for DDP training. Figure 4 shows the details of the mel-based discriminator. We combine the mel-spectrogram and log-mel-spectrogram features and then input them into a network consisting of several convolutional layers. Our motivation is that the mel-spectrogram has a strong intrinsic inductive bias, especially for sounds and music-related audio (the SOTA sounds or music classification systems are based on the log-mel-spectrogram in the literature.). Thus, we speculate that choosing a mel-spectrogram-based discriminator can better promote high-fidelity audio reconstruction. In our experiments, we use 6 different discriminators with different configurations  [^10] In our experiments, we find the mel-based discriminator brings better reconstruction performance when we train a universal neural audio codec.. Specifically, we set the hidden_dim as {64, 128, 256, 512, 512, 512} and the hop length as {32, 64, 128, 256, 512, 1024}. We train the neural audio codec model based on the Librilight and AudioSet datasets. Table 16 demonstrates that the neural codec model adopted in this work outperforms prior Encodec (encodec).
Figure 4: The overview of a single Mel-based discriminator. In practice, we will use multiple discriminators by setting different hop lengths and hidden dimensions.
Appendix F Subjective Evaluation
For TTS and VC tasks, we focus on speech quality (QMOS) and speaker similarity (SMOS). The details are as follows. For speech quality evaluation, we conduct the MOS (mean opinion score) tests and explicitly ask the raters to focus on examining the audio quality and naturalness, and ignore the differences of style (timbre, emotion, and prosody. The testers present and rate the samples, and each tester is asked to evaluate the subjective naturalness on a 1-5 Likert scale.
For speaker similarity evaluation, we ask the raters to focus on the similarity of the speaker identity (timbre) to the reference, and ignore the differences in content, grammar, or audio quality. We paired each synthesized utterance with a reference utterance to evaluate how well the synthesized speech matched that of the target speaker.
For SE and TSE tasks, we write explicit instructions to ask the rater to assess the generated speech. Refer to Figure 5 to see the details.
For SVS, we also conduct quality MOS (QMOS) and style similarity MOS (SMOS). Different from TTS’s SMOS evaluation, we explicitly instruct the raters to focus on the similarity of the style (timbre, emotion, and prosody) to the reference, and ignore the differences in content, grammar, or audio quality.
For sound and music generation tasks, we follow AudioGen (kreuk2022audiogen) and MusicGen (musicgen) to evaluate (1) overall quality (OVL), and (2) relevance to the text input (REL).
Our subjective evaluation tests are crowd-sourced and conducted by 20 native speakers via Amazon Mechanical Turk. The screenshots of instructions for testers have been shown in Figure 5. We paid about \$500 on participant compensation. A small subset of speech samples used in the test is available at https://uniaudio666.github.io/demo_UniAudio/.
Figure 5: Screenshots of subjective evaluations.